# The Dynamic Net Architecture: Learning Robust and Holistic Visual Representations Through Self-Organizing Networks

### v1

@@ 6.2 Reconstructing Subtractive Noise
Figure 6: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the “activation probabilities” (on the right) after the first time step. The activation probability is close to $1$ where a line has been observed and around $50\%$ in the middle where no line was observed but expected. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ plays a crucial role. The potential is higher at spatial locations where line features are observed and lower where they are not observed but expected. Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 6. Modeling this ambiguity enables having different internal interpretations, such as whether “it is an interrupted line that should be reconstructed” or “there are two co-linear line segments.” We find that the activation map is especially reliable during the first time step, where it is calculated based on the observed input and the network’s expectation about local patterns. Since local interactions can trigger or deactivate neurons, the activation map changes over time, reducing its interpretability.
@@
The reconstruction process unfolds across multiple sequential steps, as illustrated in Figure 7. Initially, at time step $t=0$, the net fragments represent the interrupted line. Subsequently, during the second time step ($t=1$), some of the removed pixels are reconstructed, and shortly after, at $t=2$, the entire line is represented by the net fragments. Thus, reconstruction can persist even amid increasing inhibition, given that inactive neurons receive adequate support. This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of certain neurons encourages others to follow.
6.3 Compositionality
A single DNA area models a feature hierarchy through nested net fragments of increasing size (von der Malsburg et al., 2022b). In this study, we demonstrate that these higher-level nets can be composed of local features (small net fragments), even though the underlying patterns had never been observed during training. We show this on the example of curved lines, which share only local features with the straight lines of the training set.
Figure 8: Input images of curved lines are shown on the left, the second and third columns show the corresponding feature and output activation without noise, and the fourth and fifth columns show the same feature and output activation with noise corruption (activation bias $b^{(S2)}=0.3$, inhibition coefficient $\gamma=0.6+0.1t$).
Figure 8 depicts four samples of curved lines with the corresponding neuronal activation within an area, once without noise and once with $18.5\%$ noise added to its afferent input. The cortical area is able to represent the curved lines as composite of net fragments learned from straight lines. Added unstructured clutter is efficiently removed, leading to robust input representation.
7 Discussion and Conclusions
The concepts proposed here may be compared to established neural theory. Nets are based on associative learning but differ fundamentally from Hebbian assemblies (see Appendix A) as described in (Hopfield, 1982; Cohen and Grossberg, 1983; Papadimitriou et al., 2020), which are monolithic and without inner structure, whereas nets can represent an infinity of patterns based on shared fragments and can deal with new entities based on homeomorphy between common compositional structure, thus enabling the systematicity and productivity characterizing cognitive operations (Fodor and Pylyshyn, 1988; von der Malsburg, 2023). In comparison to multi-layered perceptrons, coherent nets replace hierarchies of independent feature neurons by net fragments, share with them the potential for equivariance (shift symmetry), but in distinction to them support the construction of an invariant representation that doesn’t need to be learned object-by-object (see next section). With the transformer architecture (Vaswani et al., 2017), nets share the provision for finding non-local patterns through on-the-fly computation of lateral links.
A decisive difference between DNA and all deep-learning systems (LeCun et al., 2015; Schmidhuber, 2015; Stadelmann et al., 2019) is the mode of learning. ANNs typically rely on backpropagation of error where the goal is reaching consistency (minimal error) between system output and teaching signals. With DNA models, the goal is mutual consistency at each neuron, between the excitatory connections converging there, this consistency being measured by the probability for each connection to reach a highly active neuron when carrying a signal itself. Of all possible connectivity patterns, only a minute fraction has this property of self-consistency. This condition of self-consistency acts as a selection criterion for eligible networks and thus powerfully reduces the search space of learning.
Net fragments and the structured nets they build form a basis for the representation of hierarchical structures. They are effective for feature filtering and for preventing early commitment, which may be of fundamental importance for improving the robustness of current machine learning systems. So far, a neural code based on net fragments has merely been proposed as a theoretical concept (von der Malsburg et al., 2022b), while this study pioneers the implementation of net fragments in a concrete simulation.
Our experiments demonstrate that the proposed DNA architecture builds net fragments that effectively stabilize patterns, enhance robustness by eliminating visual clutter, and can reconstruct removed parts of known patterns. The DNA model can handle additive noise exceptionally well and eliminates up to $96\%$ of it while preserving the initial pattern, even in scenarios where up to $59\%$ of all locations are affected. Figure completion works well for line interrupts for segments of up to $3$ pixels, while for larger interruptions, the reconstruction rate decreases considerably. One reason for this is certainly due to the nature of the used 1D pattern: as it involves straight lines with a width of one pixel, typically, only a fraction of the neurons within the local 2D neighborhood of size $h^{(W2)}\times w^{(W2)}$ are active. With a larger line interrupt, the proportion of active neurons compared to neighboring neurons is vanishingly small, which makes reconstruction more difficult. It is expected that reconstruction will be much more robust with two-dimensional patterns.
We posit that this work, which focuses on the internals of a single cortical area, allows for scaling to multiple areas that will allow the implementation of more complex tasks (see next section). Nevertheless, although this work is intended to be a step towards implementing a robust learning framework and the proposed neural code exhibits promising characteristics, it still has weaknesses that should be addressed in future work. One such weakness is dependence on manual adjustment of the parameters $b$ and $\gamma$. The mechanism of signal control is biologically implausible. It should be replaced by a more elegant, theoretically justified, and self-adjusting mechanism.
8 Future Work
One important purpose of handing structure from one area to another is generalization, a prime example for which in vision is object recognition invariant to translation (and wider transformation groups). While our system only includes one “cortical” area, another area (that would correspond to the cortical inferotemporal cortex (Ito et al., 1995)) could contain invariant object models. In the context of DNA, we see invariant recognition as realized by a process in which an object-representing net in a primary sensory area $A$ forms and activates a homeomorphic net in a secondary area $B$. Two nets are homeomorphic if they contain the same feature types in the same spatial arrangement. Invariance is achieved if differently transformed mutually homeomorphic nets in $A$ activate the same net in $B$ under a homeomorphic mapping. Inter-areal homeomorphic maps are based on feature type-preserving connections, which in the brain are realized by axonal fibers [^2] With the help of several intermediate areas, the number of necessary connections can be brought down to realistic values (Anderson and van Essen, 1987; Olshausen et al., 1993; Wolfrum et al., 2008)).. In fact, connections should not only be feature type-preserving but also restricted to pairs of corresponding alternative neurons in $A$ and $B$, such that they carry not only feature information but also information on connectedness [^3] To further ease the formation of homeomorphic maps in $B$, a mechanism for mutual support between mapping-compatible connections has been proposed and successfully applied (see Section 2.3).. The generation of the net in $B$ is again the result of a two-phase process, the afferent connections from $A$ to $B$ transiently activating many neurons in $B$, most of which getting silenced in phase two, leaving only those neurons active that not only got afferent input but are also supported by lateral excitation within $B$. The net thus formed in $B$ is invariant to homeomorphic transformations in $A$. Object classification, the identification of the net in $B$ as a particular object type or recognition of individual objects needs a further area $C$ containing a fixed net that is activated based on approximate homeomorphy with the invariant net in $B$.
The result of the recognition process is the formation of a larger net composed of a net in $A$, a net in $B$, and one-to-one connections between corresponding neurons in $A$ and $B$. The self-consistency of this larger net is the basis for learning all participating connections, as all connections are stabilized (or generated) by their success in predicting activity in their target neuron.
Through inverse mapping, these stabilizing interactions can be re-imported to the previous area and thus help to discriminate the corresponding nets there from noise and clutter and bridge larger partial occlusions, which would realize the idea of predictive coding Keller and Mrsic-Flogel (2018).
An obvious first step in preparation for a larger-scale system of interacting areas is, however, the training of a single primary-sensory area with natural images to capture as net fragments all textures that occur in them with significant frequency. This project should include the extraction of such different sensory qualities as stereo depth, local motion or surface discontinuities. Net fragments and nets based on these would then have a chance to relate to intrinsic surface properties such as form.
@@ 8 Future Work
This research has been funded by the Canton of Zurich, Switzerland, through the Digitalization Initiative of the Canton of Zurich (DIZH) Fellowship project “Stability of self-organizing net fragments as inductive bias for next-generation deep learning.”
@@
The code to generate the data as well as to reproduce the results can be found after the publication of this work on Github at https://github.com/sagerpascal/dynamic-link-architecture.
@@
Appendix A Net Fragments vs. Associative Memory, Alternative Neurons
Learning of lateral connections to form net fragments shares aspects with associative memory (Hopfield, 1982; Cohen and Grossberg, 1983). There are, however, important differences. In an associative memory system, neural patterns are stored at once in their entirety with the help of synaptic plasticity, any pair of active neurons within a pattern strengthening their mutual excitatory connections. The resulting connectivity turns each stored pattern into an attractor state that can be recalled (dynamically activated) starting from initial states that come near to it. Such stored and dynamically stabilized neural patterns are often called “assemblies” (Hebb, 1949), and are discussed to this day as fundamental to the function of the brain, see for instance (Papadimitriou et al., 2020).
There are two problems with the concept of associative memory. One is rather fundamental in that a pattern is monolithic and can be recalled only as a rigid whole. This is due to the indiscriminate stickiness with which any pair of simultaneously active neurons is tied together by synaptic plasticity in the one-time storage process. Fragments, in distinction, are formed in a protracted statistical process in which connections are gradually established between neurons that are co-active with statistical significance. Due to the spatial deformation of visual patterns, this can only be the case for neurons relating to neighboring visual points. DNA models distinguish insignificant from significant connections, only the latter being permitted to grow, which they do under competition until reaching equilibrium. As a result, the system doesn’t store rigid global patterns but is able to dynamically stabilize an unlimited number of global states that are composed combinatorially out of overlapping fragments.
The other, more technical, problem of associative memory is that stored states have to be statistically independent (“orthogonal”) to avoid cross-talk between them. (In a simple example of cross-talk, a neuron $c_{1}$ gets connected at different times with neurons $c_{2},c_{3},c_{4}$, although the pattern $\{c_{1},c_{2},c_{3},c_{4}\}$ doesn’t occur with statistical significance during learning. As a result, $c_{1}$ may be activated erroneously if $\{c_{2},c_{3},c_{4}\}$ is active as part of a larger pattern.) Sparsity (according to which a very small fraction of all neurons are active within any given pattern) has been proposed as a general means to avoid cross-talk and to ensure statistical independence of stored patterns (French, 1999).
DNA models are freed from this constraint by the introduction of “alternative neurons,” neurons that are activated by the same input feature but are free to learn different lateral connections to other neurons. (In our example, neuron $c_{1}$ would be replaced by a set $\{c_{1},c_{1^{\prime}},c_{1^{\prime\prime}}\}$ of alternative neurons, which could then independently learn the connections $c_{1}-c_{2}$, $c_{1^{\prime}}-c_{3}$ and $c_{1^{\prime\prime}}-c_{4}$, thus avoiding the above cross-talk.) In the biological case, there is ample room for alternative neurons. There are, for instance, a hundred times more neurons in the primary visual cortex compared to the number of fibers coming out of the retina (Leuba and Kraftsik, 1994).
Appendix B Weight Initialization
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are copied into the $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative cells.
The lateral connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$. When Hebbian plasticity is employed on the weight matrix $\boldsymbol{W}^{(L)}$, these connections ought to grow (as neurons tend to keep their activation state between time steps), but due to the constraint of keeping all connection weights within the range $(0,1)$ (refer to equation 8), they stay put at $1$.
Appendix C Hebbian Plasticity
@@
where $\alpha$ represents the learning rate, and $\boldsymbol{\rho}_{avg}$ denotes the average correlation between input and output. This average is determined for a given activation over all instances of a given pair of neurons $c,c^{\prime}$ in different positions of the image domain. To compute $\boldsymbol{\rho}_{avg}$, we count the number of instances in which both neurons are on, subtract from it the number of instances where one neuron is on, and the other is off, and divide that sum by the number of all positive and negative instances. The actual modification eq. (8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
The network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant. Consequently, the same lateral connection is applied across multiple output and input neurons. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects. If the resulting average is positive, the corresponding connection increases by $\alpha$ as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced by $\alpha$.
Appendix D Normalization
We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$. These steps are functionally motivated and not by reference to biology.
@@ Appendix D Normalization
$\boldsymbol{a}^{\prime\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
@@
According to this formula, $\boldsymbol{a}^{\prime\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$. (In our experiments we found the value $\lambda=1.3\cdot\frac{h^{(W2)}+w^{(W2)}}{2}$ to work well). This saturation helps to mitigate imbalances between network fragments with differing numbers of participating neurons (as with our stimuli, line segments, where many more feature neurons activate around the end-points than in the middle of segments).
In the next step, we introduce the normalization
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
@@
This normalization confines activations to the interval $(0,1)$, where neurons receiving maximal support are mapped to $1$, while those with lesser support are mapped to correspondingly smaller values. This normalization relative to the maximal value that itself is growing during learning instead of relative to a constant parameter like $\lambda$ is important to give neurons initially (when lateral connections are still small) a chance to fire and later, when lateral links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
Appendix E Alternative Neuron Selection

### v2

@@ 5.3 Reconstructing Figures
When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected. Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 7B. Modeling this ambiguity enables having different internal interpretations, such as whether ‘it is an interrupted line that should be reconstructed’ or ‘there are two co-linear line segments’ and selecting one of them based on the threshold $b^{(S2)}$.
@@
The reconstruction process unfolds across multiple sequential steps, as illustrated in Figure 8. Initially, at time step $t=0$, the net fragments represent the interrupted line. Subsequently, during the second time step ($t=1$), one of the removed pixels is reconstructed, and shortly after, at $t=2$, the entire line is represented. Thus, reconstruction can persist even amid increasing inhibition, given that inactive neurons receive adequate support. This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of neurons encourages others to follow.
5.4 Comparison to an Autoencoder
Figure 9: Comparison of the outputs generated by the proposed CNA model and an autoencoder across different tasks. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows $1$ to $3$ illustrate how both models encode patterns that were not present in the training dataset. Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
Figure 9 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix H). While the autoencoder successfully learns to accurately represent line structures in the training data - achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and decoder reconstruction - it struggles with generalization to data not observed during training. For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
Rows $4$ and $5$ highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced. However, as the noise level increases, its performance decreases significantly. The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations. Metrics on noise reduction, precision, and recall, presented in Appendix I, further substantiate the CNA model’s superior performance.
In row $6$, the models’ capabilities in reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This is confirmed by the metrics in Appendix I: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of more noise at other locations, leading to lower overall precision and recall.
6 Discussion and Outlook
The CNA presented here establishes a novel approach to construct coherent nets as neural representations of visual input, enabling arrays of elements to be treated as unified gestalts through the composability of previously learned sub-patterns (net fragments). While the concept of such net construction has been proposed theoretically (von der Malsburg et al., 2022), this work marks the first successful implementation and simulation of the idea.
Our approach offers distinct advantages over existing neural coding frameworks, particularly in terms of robustness and the ability to represent out-of-domain data. Using nets of neurons to represent sensory signals allows filtering out noise and figure completion, while using net fragments allows representing unseen objects through composing small feature patches. In contrast to past work that relied on manually constructed filter functions (e.g., Geman and Geman, 1984), the filter functions, nets in our scheme, are learned from input statistics resp. constructed on the spot, so that they will eventually be able to cover the whole gamut of patterns encountered in natural images.
An essential next step would be the construction of a system that is trained on natural images, that contains at each position of the visual field a complement of elementary feature detector neurons able to accurately represent the pixels of the input (as in Olshausen and Field, 1996), and whose recurrent connections have sufficient lateral extent to capture the most important signal correlations present in natural images. Such systems can then be expected to develop net fragments that capture and represent all textures encountered in natural images on that length scale, thus realizing the ambition expressed in Figure 1.
A first and important such operation is abstracting over position, size, and in-plane orientation of an object’s image in the input area (primary visual cortex) by the construction of an invariant net representation of the object in a separate area (which would correspond to inferotemporal cortex (Ito et al., 1995)). As hinted at in the Introduction, this can be based on structure-sensitive, ‘homeomorphic’, mappings, as proposed in von der Malsburg (1981); Anderson and Van Essen (1987); Olshausen et al. (1995); Wiskott et al. (1997); Wolfrum et al. (2008); Memisevic and Hinton (2010) or Arathorn et al. (2013). A mapping is homeomorphic if it preserves neural feature type as well as connectivity. Object classification or recognition of individual objects can then be based on a further homeomorphic mapping to an area containing a more permanent schematic representation of the object class or a sample of the object to be recognized, as realized in Wolfrum et al. (2008).
The introduction of nets as a neural code and the inclusion of alternative neurons suggest a natural basis for the mapping process by requiring inter-areal connections to be not only specific to feature type but also to alternative neuron identity. As this identity contains information on net connectivity, this specificity can implement the homeomorphy constraint of inter-areal mappings in a natural way and overcome the limitations in existing systems (Wolfrum et al., 2008).
Once the mechanism of homeomorphic mapping has been established, it is available for arbitrary novel objects, thus forming the basis for invariant object recognition at first sight (as observed in humans in Biederman and Cooper (1991)). This capability stands in stark contrast to current ANN technology, which requires invariance to be trained individually for each object, highlighting the broader generalization potential of our approach.
The simple mechanism of net formation that we model here is a first but significant step in the direction of a data architecture grounded in nets connecting sparse sets of neurons within and between areas. Nets are singled out by the properties of cooperativity and inner consistency: all neurons of an active net experience the high-order coincidence of converging excitatory signals, realizing predictive coding at each neuron. Given high sparsity of neural firing and high sparsity of connectivity, satisfying the requirement of consistency is highly non-trivial and can only be established by a learning process in which single neurons acquire converging connections coming from sets of neurons that fire simultaneously again and again, as realized in our model.
In summary, the CNA demonstrates the feasibility and potential of a principled and compositional approach to neural coding. Its ability to autonomously construct robust representations, handle out-of-domain data, and generalize to novel stimuli underscores its promise as a foundation for more advanced systems in the future, especially when extended with multiple areas.
@@ 6 Discussion and Outlook
This research has been funded by the Canton of Zurich, Switzerland, through the Digitalization Initiative of the Canton of Zurich (DIZH) Fellowship project ‘Stability of self-organizing net fragments as inductive bias for next-generation deep learning.’
@@
The code to generate the data as well as to reproduce the results can be found after the publication of this work on Github at https://github.com/sagerpascal/cna.
@@
Appendix A Weight Initialization
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are copied into the $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative neurons.
The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$. When Hebbian plasticity is employed on the weight matrix $\boldsymbol{W}^{(L)}$, these connections ought to grow (as neurons tend to keep their activation state between time steps), but due to the constraint of keeping all connection weights within the range $(0,1)$ (refer to equation 8), they stay put at $1$.
Appendix B Hebbian Plasticity
@@
where $\alpha$ represents the learning rate. Note that the network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant. Consequently, the same recurrent connection is applied across multiple output and input neurons. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects. If the resulting average is positive, the corresponding connection increases by $\alpha$ as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced by $\alpha$.
The actual modification eq. (8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
Appendix C Normalization
We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$.
@@ Appendix C Normalization
$\boldsymbol{a}^{\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
@@
According to this formula, $\boldsymbol{a}^{\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$. (In our experiments we found the value $\lambda=1.3\cdot\frac{h^{(W2)}+w^{(W2)}}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
@@
This normalization confines activations to the interval $(0,1)$, where neurons receiving maximal support are mapped to $1$, while those with lesser support are mapped to correspondingly smaller values. This normalization relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later, when recurrent links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
Appendix D Alternative Neuron Selection